An object of volume 4x10^4m^3 and density 6000kg/m^3 is immersed inside a fluid of density 5000kg/m^3. The force exerted by the fluid on the object is 0.7N. Calculate the density of the fluid
1
Expert's answer
2021-09-14T09:36:16-0400

According to this principle, the buoyant force on an object equals the weight of the fluid it displaces. In equation form, Archimedes’ principle is


Fwfl - mobjectg


where FB is the buoyant force and wfl is the weight of the fluid displaced by the object.


"F_B =0.7\,N
\\F_B=w_{fl}-m_{object}\cdot g=\rho_{fl}\cdot g\cdot V_{object}-V_{object}\cdot \rho_{object} \cdot g
\\ \text{ }
\\ \implies \rho_{fl}=\dfrac{F_B+V_{object}\cdot \rho_{object} \cdot g}{g\cdot V_{object}}=\dfrac{F_B}{g\cdot V_{object}}+\rho_{object}
\\ \text{ }
\\ \implies \rho_{fl}=\dfrac{0.7\,N}{(9.81\,{m/s^2})(4\times10^{-4}\,m^3)}+6000\, \dfrac{kg}{m^3}
\\ \text{ }
\\ \implies \rho_{fl}=6178.39\dfrac{kg}{m^3}"


In conclusion, the density of the fluid is about 6178.39 kg/m3.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.